# Beyond Chain-of-Thought: A Survey of Chain-of-X Paradigms for LLMs

### v1

@@ 4.3 Multi-Step Reasoning
Reasoning has been a widely studied topic, particularly multi-step reasoning tasks that demand a robust understanding of context and logic Wei et al. (2022). The sequential nature of CoX methods makes them ideally suited for this task. For instance, Chain-of-Knowledgeb Wang et al. (2023c) elicits explicit knowledge evidence at each step, thereby improving LLMs’ performance in various reasoning tasks. Meanwhile, Chain-of-Feedback Ahn and Shin (2024) revises initial incorrect reasoning steps by breaking them down into smaller, individual tasks for more grounded reasoning. Other specialized reasoning tasks include rule-based reasoning Servantez et al. (2024), database reasoning Hu et al. (2023a), legal reasoning Kuppa et al. (2023), user behavioral reasoning Do et al. (2023); Han et al. (2024), structure and graph reasoning Zeng et al. (2024); Luo et al. (2024); Xia et al. (2024d), as well as reasoning for text summarization Adams et al. (2023); Bao et al. (2024) and machine translation Lu et al. (2023). Through these varied applications, CoX methods demonstrate their ability to decompose complex tasks into manageable steps, enhancing LLMs’ ability to process and analyze information effectively.
@@ 4.4 Instruction Following
Instruction following, as a celebrated abilities of LLMs, enables humans to provide explicit guidance for various tasks Zhang et al. (2023b). The evolution of CoX methods has also led to various approaches for enhancing this feature. Chain-of-Task Li et al. (2024b), for instance, offers a structured method where each instruction consists of intermediate atomic tasks, specifically curated to fine-tune an e-commerce LLM’s responses to better meet customer needs. Extending this concept, Chain-of-Instructions Hayati et al. (2024) introduces a compositional approach where each output feeds directly into the next, creating a continuous loop of task-specific tuning that refines the LLM’s task handling progressively. For applications in speech generation, Chain-of-Modality Zhang et al. (2023a) constructs fine-tuning sequences with concatenated textual and speech instructions. Furthermore, Chain-of-LoRA Qiu et al. (2024) uses LoRA networks to specialize instruction handling, optimizing performances across varied tasks by tailoring instruction tuning processes to each LoRA. These developments underscore how CoX methods can improve the instruction-following capabilities of LLMs, enabling them to understand and execute tasks with more clarity.
@@ 4.5 LLMs as Agents
With the strong planning abilities, LLMs have been utilized as agents across a wide range of tasks Xi et al. (2023b). CoX methods have been explored to further boost the planning abilities of LLM-based agents. In this vein, Chain-of-Actiona Zhan and Zhang (2023) and Chain-of-ActionThought Zhang et al. (2024c) utilize a series of planned actions to guide agent decision-making, ensuring each step is informed by the previous one. While in games like StarCraft II, Chain-of-Summarization Ma et al. (2023) employs LLMs to summarize past observations to suggest future strategies. Chain-of-3DThought Yamada et al. (2024) further utilizes LLM agents to compose objects in images through trial and error within a 3D simulation environment. LLMs also serve as planners in human-scene interaction tasks with Chain-of-Contacts Xiao et al. (2024a), and in tool using tasks with Chain-of-Abstraction Gao et al. (2024). CoX methods have also been applied to multi-agent settings, such as Chain-of-Experts Xiao et al. (2024b) and Chain-of-Discussion Tao et al. (2024). These methods highlights the integration of CoX methods in enhancing the multi-dimensional abilities of LLMs as autonomous and collaborative agents.
@@ 4.6 Evaluation Tools
Evaluating LLMs has become increasingly challenging as they grow more sophisticated Chang et al. (2023), making CoX methods a valuable asset for evaluation purposes. Chain-of-Utterances prompting Bhardwaj and Poria (2023) is a prime example of how CoX methods can illuminate specific areas of concern, such as safety issues in scenarios where LLMs interact with potentially harmful models. This method uncovers vulnerable conversations that could lead to LLM jailbreaking, providing essential insights into the robustness of LLM safety. Besides, Chain-of-Feedback Ahn and Shin (2024) method has demonstrated the impact of prompts on LLM performance. By repeatedly providing LLMs with non-informative prompts like “make another attempt”, researchers have observed a gradual decrease in the quality of responses. In visual reasoning, the Chain-of-Images Meng et al. (2023) introduces a benchmark involving a sequence of images designed to progressively assess the LLM’s reasoning ability. It provides a robust tool for measuring the model’s capabilities in interpreting visual data. These CoX methods underscore the importance of nuanced evaluation for more thorough assessments of LLMs.
@@ 5 Future Directions
Existing works generally focus on improving task-specific generative results. However, understanding and explaining the underlying mechanisms of LLM reasoning is also essential in realistic scenarios. For example, Wang et al. (2023d) show that LLMs may skip rational steps when generating final results. Wang et al. (2023a) observe a performance gain from CoT even with invalid rationales. These observations indicate the value of a causal analysis on how intermediate steps truly affect the final results.
@@
A chain leading to the final node of generation often requires multiple sequential inference steps, which are computationally heavy and time-consuming, especially with LLMs. It would be interesting if future research could reduce the length of CoX chains while maintaining the quality of generation. For example, it would be worth studying whether the intermediate nodes of CoX could be executed in parallel or jointly within a single inference step.
@@ Appendix A Taxonomies of Nodes and Tasks
[CoUtterances Bhardwaj and Poria (2023), CoImages Meng et al. (2023), CoExplanation Huang et al. (2023), CoAspects Gong and Mao (2023), CoFeedback Ahn and Shin (2024), text width=10.9cm, fill=lightblue]
@@
[CoActiona Zhan and Zhang (2023), CoSymbol Hu et al. (2023b), CoSummarization Ma et al. (2023), Co3DThought Yamada et al. (2024), CoActionThought Zhang et al. (2024c), CoExperts Xiao et al. (2024b), CoDiscussion Tao et al. (2024), CoAbstraction Gao et al. (2024), CoContacts Xiao et al. (2024a), text width=10.9cm, fill=lightblue]
@@
[CoTask Li et al. (2024b), CoImagination Zhou et al. (2024), CoInstructions Hayati et al. (2024), CoModality Zhang et al. (2023a), CoLoRAb Qiu et al. (2024), text width=10.9cm, fill=lightblue]
@@
[CoT Wei et al. (2022), CoDensity Adams et al. (2023), CoKnowledgeb Wang et al. (2023c), CoMemory Hu et al. (2023a), CoOpinion Do et al. (2023), CoReference Kuppa et al. (2023), CoQuery Xu et al. (2023), CoInteraction Han et al. (2024), CoLogic Servantez et al. (2024), CoFeedback Ahn and Shin (2024), CoLayer Zeng et al. (2024), CoEvent Bao et al. (2024), text width=10.9cm, fill=lightblue]
@@
[CoUtterances Bhardwaj and Poria (2023), CoHindsight Liu et al. (2024a), text width=9.0cm, fill=lightblue]
[Hallucination Reduction, text width=1.3cm, l sep = 2mm
[CoNLI Lei et al. (2023), CoVerification Dhuliawala et al. (2023), CoKnowledgea Li et al. (2024a), CoNote Yu et al. (2023a), CoQuestion Huang et al. (2024), CoActionb Pan et al. (2024), text width=9.0cm, fill=lightblue]
@@
[CoInformation Zhang et al. (2024a), CoModality Zhang et al. (2023a), text width=9.0cm, fill=lightblue]
@@
[CoRepair Wang et al. (2023b), CoCode Li et al. (2023a), CoSimulation La Malfa et al. (2024), CoSelfRevisions Le et al. (2024), text width=9.0cm, fill=lightblue]
@@
[CoCommand Zha et al. (2023), CoTable Wang et al. (2024), text width=9.0cm, fill=lightblue]
@@
[CoInstructEditing Zhang et al. (2023d), CoLook Xi et al. (2023a), CoQA Kim et al. (2024), CoSpot Liu et al. (2024b), CoReasoning Uehara et al. (2024), CoManipulation Qi et al. (2024), text width=9.0cm, fill=lightblue]
@@
[CoExperts Xiao et al. (2024b), CoDiscussion Tao et al. (2024), CoLoRAa Xia et al. (2024a), CoLoRAb Qiu et al. (2024), text width=10.9cm, fill=lightblue]
@@
[Self-Refine, text width=1.3cm, l sep = 2mm
[CoNLI Lei et al. (2023), CoVerification Dhuliawala et al. (2023), CoDensity Adams et al. (2023), CoSelfRevisions Le et al. (2024), CoFeedback Ahn and Shin (2024), text width=9.0cm, fill=lightblue]
[External
Feedback, text width=1.3cm, l sep = 2mm
[Co3DThought Yamada et al. (2024), CoRepair Wang et al. (2023b), CoHindsight Liu et al. (2024a), text width=9.0cm, fill=lightblue]
@@
[CoEmpathy Lee et al. (2023b), CoReference Kuppa et al. (2023), CoDictionary Lu et al. (2023), CoMemory Hu et al. (2023a), CoAbstraction Gao et al. (2024), text width=9.0cm, fill=lightblue]
@@
[CoQuery Xu et al. (2023), CoNote Yu et al. (2023a), CoKnowledgea Li et al. (2024a), CoQuestion Huang et al. (2024), CoActionb Pan et al. (2024), text width=9.0cm, fill=lightblue]
@@
[CoOpinion Do et al. (2023), CoActiona Zhan and Zhang (2023), CoSummarization Ma et al. (2023), CoLayer Zeng et al. (2024), CoHistorya Luo et al. (2024), CoHistoryb Xia et al. (2024d), text width=9.0cm, fill=lightblue]
@@
[CoSymbol Hu et al. (2023b), CoKnowledgeb Wang et al. (2023c), CoManipulation Qi et al. (2024), CoSimulation La Malfa et al. (2024), CoSpot Liu et al. (2024b), CoReasoning Uehara et al. (2024), text width=9.0cm, fill=lightblue]
@@
[CoT Wei et al. (2022), CoCode Li et al. (2023a), CoTable Wang et al. (2024), CoLogic Servantez et al. (2024), CoEvent Bao et al. (2024), CoInteraction Han et al. (2024), CoActionb Pan et al. (2024), CoInformation Zhang et al. (2024a), text width=9.0cm, fill=lightblue]

### v3

@@ 4.3 Multi-Step Reasoning
Multi-step reasoning typically demands a robust understanding of context and logic Wei et al. (2022). These tasks require breaking down complex problems into a series of smaller, interconnected steps, building upon each step to reach a logical conclusion. The sequential nature of CoX makes it ideally suited for these tasks, including rule-based reasoning Servantez et al. (2024), database reasoning Hu et al. (2023a), legal reasoning Kuppa et al. (2023), user behavior reasoning Do et al. (2023); Han et al. (2024), graph reasoning Zeng et al. (2024); Luo et al. (2024); Xia et al. (2024e), as well as reasoning for summarization Adams et al. (2023); Bao et al. (2024) and machine translation Lu et al. (2023). These varied applications demonstrate CoX’s advantages in enhancing LLMs’ ability to process information more effectively.
@@ 4.4 Instruction Following
Instruction following has been a celebrated ability of LLMs Zhang et al. (2023b). The evolution of CoX methods has also led to various approaches for enhancing this feature. A notable line of works, such as Chain-of-Task Li et al. (2024b), LogiCoT Liu et al. (2023) and Chain-of-Imagination Zhou et al. (2024), construct sequences of instructions for prompting or instruction tuning to handle complex tasks that require explicit step-wise guidance. While training a single LLM to follow different instructions can be costly, Chain-of-LoRA Qiu et al. (2024) adopts a series of LoRA networks to specialize in instruction handling. After identifying an instruction type, Chain-of-LoRA applies task-specific LoRA networks to the base LLM to accomplish the respective tasks.
@@ 4.5 LLMs as Agents
The planning abilities have made LLMs strong agents across a wide range of tasks Xi et al. (2023b). CoX methods have been explored to further boost the planning abilities of LLM-based agents. In this vein, Chain-of-Actiona Zhang and Zhang (2024) and Chain-of-ActionThought Zhang et al. (2024b) utilize a series of planned actions to guide the decision-making of agents, ensuring each step is informed by the previous actions. As discussed previously, LLM-based agents can also be augmented with historical data, e.g., Chain-of-Summarization Ma et al. (2023), and external model feedback, e.g., Chain-of-3DThought Yamada et al. (2024). LLMs also serve as planners in human-scene interaction tasks with Chain-of-Contacts Xiao et al. (2024a), and in tool using with Chain-of-Abstraction Gao et al. (2024). Chain-of-Models also naturally involves multi-agent settings such as Chain-of-Discussion Tao et al. (2024). These methods highlight the integration of CoX in enhancing the multi-dimensional abilities of LLMs as autonomous and collaborative agents.
@@ 4.6 Evaluation Tools
Evaluating LLMs has become increasingly challenging as they grow more sophisticated Chang et al. (2023), making CoX methods a valuable asset for evaluation purposes. Chain-of-Utterances prompting Bhardwaj and Poria (2023) exposes safety issues where LLMs interact with potentially harmful models. BadChain Xiang et al. (2024) also reveals vulnerabilities of LLMs outputting unintended malicious content under backdoor attack when employing CoT prompting. Chain-of-Feedback Ahn and Shin (2024) conducts another evaluation, demonstrating that by repeatedly providing LLMs with non-informative prompts like “make another attempt”, the quality of responses gradually decreases. These methods underscore the importance of nuanced evaluations of LLMs.
@@ 5 Future Directions
Existing works generally focus on improving task-specific generative results. However, understanding and explaining the underlying mechanisms of LLM reasoning is also essential in realistic scenarios. For example, Wang et al. (2023f) show that LLMs may skip rational steps when generating final results. Wang et al. (2023a) observe a performance gain from CoT even with invalid rationales. These observations indicate the value of a causal analysis on how intermediate steps truly affect the final results.
@@
A chain leading to the final node of generation often requires multiple sequential inference steps, which are computationally heavy and time-consuming, especially with LLMs. Future research may explore reducing the length of CoX chains while maintaining the quality of generation. It would also be worth studying whether the intermediate nodes of CoX could be executed in parallel or jointly within a single inference step.
@@ 6 Conclusion
Limitations
This survey presents a comprehensive overview of Chain-of-X methods. We have made our best efforts to collect studies leveraging the CoX concept, regardless of whether they are explicitly named as such. However, with the rapidly growing number of works in the field, there is still a chance that some have been missed. We welcome suggestions from the research community and will continue our efforts to survey and update the collected works.
@@ Appendix A Taxonomies of Nodes and Tasks
[CoUtterances Bhardwaj and Poria (2023), CoImages Meng et al. (2023), CoExplanation Huang et al. (2023), CoAspects Gong and Mao (2023), BadChain Xiang et al. (2024), CoFeedback Ahn and Shin (2024), text width=10.9cm, fill=lightblue]
@@
[CoActiona Zhang and Zhang (2024), CoSymbol Hu et al. (2023b), CoSummarization Ma et al. (2023), Co3DThought Yamada et al. (2024), CoActionThought Zhang et al. (2024b), CoExperts Xiao et al. (2024b), CoDiscussion Tao et al. (2024), CoAbstraction Gao et al. (2024), CoContacts Xiao et al. (2024a), text width=10.9cm, fill=lightblue]
@@
[CoTask Li et al. (2024b), LogiCoT Liu et al. (2023), CoModality Zhang et al. (2023a), CoImagination Zhou et al. (2024), CoInstructions Hayati et al. (2024), CoLoRAb Qiu et al. (2024), text width=10.9cm, fill=lightblue]
@@
[CoT Wei et al. (2022), CoDensity Adams et al. (2023), CoKnowledgeb Wang et al. (2023d), CoMemory Hu et al. (2023a), CoOpinion Do et al. (2023), CoReference Kuppa et al. (2023), CoQuery Xu et al. (2024), CoInteraction Han et al. (2024), CoLogic Servantez et al. (2024), CoFeedback Ahn and Shin (2024), CoLayer Zeng et al. (2024), CoEvent Bao et al. (2024), text width=10.9cm, fill=lightblue]
@@
[CoUtterances Bhardwaj and Poria (2023), CoHindsight Liu et al. (2024a), AlignCoT Liu et al. (2024b), text width=9.0cm, fill=lightblue]
[Hallucination, text width=1.3cm, l sep = 2mm
[CoNLI Lei et al. (2023), CoVerification Dhuliawala et al. (2023), CoKnowledgea Li et al. (2024a), CoNote Yu et al. (2023a), KDCoT Wang et al. (2023e), CoQuestion Huang et al. (2024), CoActionb Pan et al. (2024), text width=9.0cm, fill=lightblue]
@@
[CoModality Zhang et al. (2023a), CoInformation Zhang et al. (2024a), text width=9.0cm, fill=lightblue]
@@
[CoRepair Wang et al. (2023b), CoCode Li et al. (2023a), PoT Chen et al. (2023a), CoSimulation La Malfa et al. (2024), CoSelfRevisions Le et al. (2024), text width=9.0cm, fill=lightblue]
@@
[CoCommand Zha et al. (2023), CoTable Wang et al. (2024b), TabCoT Ziqi and Lu (2023), text width=9.0cm, fill=lightblue]
@@
[CoInstructEditing Zhang et al. (2023d), CoLook Xi et al. (2023a), CCoT Mitra et al. (2024), MMCoT Zhang et al. (2023e), DDCoT Zheng et al. (2023), VisualCoT Shao et al. (2024), KAMCoT Mondal et al. (2024), CoQA Kim et al. (2024), CoSpot Liu et al. (2024c), CoReasoning Uehara et al. (2024), CoManipulation Qi et al. (2024), text width=9.0cm, fill=lightblue]
@@
[CoExperts Xiao et al. (2024b), CoDiscussion Tao et al. (2024),
ChatEval Chan et al. (2024), CoLoRAa Xia et al. (2024a), CoLoRAb Qiu et al. (2024), text width=10.9cm, fill=lightblue]
@@
[Self Feedback, text width=1.3cm, l sep = 2mm
[CoNLI Lei et al. (2023), CoVerification Dhuliawala et al. (2023), VerifyCoT Ling et al. (2023), Self-Refine Madaan et al. (2023), CoDensity Adams et al. (2023), CoSelfRevisions Le et al. (2024), CoFeedback Ahn and Shin (2024), text width=9.0cm, fill=lightblue]
[Ext. Feedback, text width=1.3cm, l sep = 2mm
[CoRepair Wang et al. (2023b), Co3DThought Yamada et al. (2024), CoHindsight Liu et al. (2024a), text width=9.0cm, fill=lightblue]
@@
[CoEmpathy Lee et al. (2023b), CoReference Kuppa et al. (2023), CoDictionary Lu et al. (2023), CoMemory Hu et al. (2023a), text width=9.0cm, fill=lightblue]
[Tools, text width=1.3cm, l sep = 2mm
[ChatCoT Chen et al. (2023b), MultiToolCoT Inaba et al. (2023), CoAbstraction Gao et al. (2024), text width=9.0cm, fill=lightblue]
@@
[CoQuery Xu et al. (2024), CoNote Yu et al. (2023a), CoKnowledgea Li et al. (2024a), Verify-and-Edit Zhao et al. (2023), ReAct Yao et al. (2023), Self-Ask Press et al. (2023), IRCoT Trivedi et al. (2023), CoQuestion Huang et al. (2024), CoActionb Pan et al. (2024), RAT Wang et al. (2024a), GraphCoT Jin et al. (2024), ToG Sun et al. (2024), text width=9.0cm, fill=lightblue]
@@
[CoOpinion Do et al. (2023), PsyCoT Yang et al. (2023), CoActiona Zhang and Zhang (2024), CoSummarization Ma et al. (2023), CoLayer Zeng et al. (2024), CoHistorya Luo et al. (2024), CoHistoryb Xia et al. (2024e), text width=9.0cm, fill=lightblue]
@@
[CoSymbol Hu et al. (2023b), CoKnowledgeb Wang et al. (2023d), CueCoT Wang et al. (2023c), CCoT Mitra et al. (2024), CoManipulation Qi et al. (2024), CoSimulation La Malfa et al. (2024), CoSpot Liu et al. (2024c), CoReasoning Uehara et al. (2024), text width=9.0cm, fill=lightblue]
@@
[CoT Wei et al. (2022), Least-to-Most Zhou et al. (2023), CoCode Li et al. (2023a), CoTable Wang et al. (2024b), CoLogic Servantez et al. (2024), CoEvent Bao et al. (2024), CoInteraction Han et al. (2024), CoActionb Pan et al. (2024), CoInformation Zhang et al. (2024a), text width=9.0cm, fill=lightblue]